OpenStudy (anonymous):

simplify 5/(x^2 - xy) - 5(y^2-xy) give detailed explanation plz I cant understand this!

OpenStudy (anonymous):

Wait is that a 5 divided by everything or a fraction subtracted from a product?

OpenStudy (anonymous):

no hold on, I didn't write it right it's 5/(x^2-xy) - 5/(y^2 -xy) it's subtracting 2 diff fractions.

OpenStudy (anonymous):

5/x(x - y) + 5/y( x- y) = 5/(x-y) [1/x -1/y] = 5/(x-y) [(y-x)/xy] -5/xy